# Function-Correcting Partition codes

### v1

@@ Abstract
We introduce function-correcting partition codes (FCPCs) that are a natural generalization of function-correcting codes (FCCs). A $t$-error function-correcting partition code is an $(\mathcal{P},t)$-encoding defined directly on a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{k}$. For a partition $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{E}\}$ a systematic mapping $\mathcal{C}_{\mathcal{P}}:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{k+r}$ is called a $(\mathcal{P},t)$-encoding if for all $u\in P_{i}$ and $v\in P_{j}$ with $i\neq j$, $d\big(\mathcal{C}_{\mathcal{P}}(u),\mathcal{C}_{\mathcal{P}}(v)\big)\geq 2t+1.$ We show that any $t$-error correcting code for a function $f$, denoted by $(f,t)$-FCC is exactly an FCPC with respect to the domain partition induced by $f$, which makes these codes a natural generalization of FCCs. We use the join of domain partitions to construct a single code that protects multiple functions simultaneously, with application for scenarios where a receiver needs to protect more than one function value or in a broadcast scenario where the different receivers need protection for different function values. We define the notion of partition redundancy gain and partition rate gain to measure the bandwidth saved by using a single FCPC for multiple functions instead of constructing separate FCCs for each function. We derive general lower and upper bounds on the redundancy of such FCPCs and illustrate the achievable gains through examples. We specialize this to linear functions via coset partition of the intersection of their kernels. Then, we associate a partition graph to any given partition of $\mathbb{F}_{q}^{k}$, and show that the existence of a suitable clique in this graph yields a set of representative information vectors that achieves the optimal redundancy. Using the existence of a full-size clique in the weight partition and support partition, we obtain lower and upper bounds on the optimal redundancy of FCPCs for these partitions. Finally, we introduce the notion of a block-preserving contraction for a partition, which helps reduce the problem of finding optimal redundancy for an FCPC. We further show that such a contraction exists for all weight-based partitions. We observe that FCPCs naturally provide a form of partial privacy, in the sense that only the domain partition of the function needs to be revealed to the transmitter.
@@ I Introduction
The theory of function-correcting codes (FCCs) has so far been developed primarily for the standard one-to-one communication setting, where a single receiver is interested in obtaining the value of a particular function of the transmitted message vector. In this work, we move beyond that setting and consider a more general and practically relevant scenario. A transmitter broadcasts a message vector over a shared and error-prone link to multiple receivers, and each receiver is interested in a different function or attribute of the same message vector. Since these receivers may require protection against errors for their respective functions, the natural question is whether one can design a single function-correcting code that simultaneously protects all the desired functions from up to $t$ errors.
A straightforward approach is to encode the entire message vector using a standard error-correcting code, thereby enabling complete recovery at all receivers. This trivially protects all functions but may introduce significantly more redundancy than necessary. Given that FCCs are typically studied in systematic form, another equally straightforward approach is to take the redundancies required for the different FCCs corresponding to the individual functions, append all of them to the message vector, and transmit the resulting longer vector. In this way, each receiver can simply use the redundancy meant for its own function to decode its desired value. However, this solution also incurs redundancy equal to the sum of all the individual FCC redundancies, which may be unnecessarily large. If, instead, one can find a unified FCC that protects all the functions with strictly smaller redundancy, it leads to substantial bandwidth savings compared to both of the above approaches: (1) using an error-correcting code that protects the entire message vector, and (2) using the sum of the redundancies of the separate FCCs designed for each function individually.
The aim is to construct codes that perform this task as required. Instead of working with the function values directly, we define codes based on partitions of the domain and refer to them as function-correcting partition codes (FCPCs). The setup is illustrated in Fig. 1. Since every function naturally partitions its domain into preimage sets, these FCPCs can be used as function-correcting codes. In fact, each function induces a natural partition of the domain into its preimage sets, and a function-correcting code is exactly the function-correcting partition code defined with respect to this induced partition. This makes FCPCs a direct generalization of function-correcting codes.
@@
Further, we refer to the savings, arising from using a single code for multiple functions instead of using separate FCCs for each function, as the partition redundancy gain and partition rate gain.
There is another benefit of using FCPCs in a network setting, as they provide a degree of partial privacy regarding the function being computed, which we refer to as function class privacy. Here, a function class consists of all functions on the domain that induce the same domain partition. In FCCs, both the receiver and the transmitter must know the exact function. In contrast, an FCPC requires the transmitter to know only the domain partition induced by the function that the receiver wishes to compute. Since many different functions can belong to the same function class and induce the same domain partition, the transmitter cannot infer the exact function from this information alone. Further discussion of this notion is given in Section VI.
@@ I-A Related works
The notion of function-correcting codes was formally introduced in [1], which laid the foundational framework for studying error correction when the objective is to reliably recover a function of the transmitted message rather than the message itself. In that work, the authors defined the key parameters of FCCs, established fundamental redundancy bounds, and investigated structural properties of such codes. They also provided explicit constructions for several important classes of functions, including locally binary functions, the Hamming weight function, and the Hamming weight distribution function. For some of these cases, the proposed constructions were shown to be optimal by achieving the corresponding lower bounds on redundancy. The initial development focused on communication over binary symmetric channels.
Subsequent research has extended this framework to more general channel models. In [2], FCCs were studied in the context of symbol-pair read channels over binary fields, and this line of work was further generalized to $b$-symbol read channels in [3].
In [4], the authors leveraged the fact that the kernel of a linear function forms a subspace of $\mathbb{F}_{q}^{k}$, and used the associated coset partition to construct FCCs for linear functions. They also derived a general lower bound on the redundancy of FCCs, which specializes to a classical lower bound for error-correcting codes in the case of bijective functions, since FCCs reduce to ECCs in this setting. The tightness of this bound was established for certain parameters. The work [5] further refined the study of FCCs for the Hamming weight and Hamming weight distribution functions, providing improved bounds and constructions tailored to these specific functions.
@@
Another line of work studies function-correcting codes under alternative distance measures and channel models. In [9, 10], FCCs are investigated in the Lee metric, where Plotkin-type bounds are derived for general functions and explicit constructions are provided for several function classes, with optimality established in certain cases. The work [11] considers FCCs with homogeneous distance, developing redundancy bounds and constructions adapted to this setting. In addition, Plotkin-type bounds for FCCs over $b$-symbol read channels are studied in [12].
More recently, a related direction was explored in [13], where the authors study function-correcting codes with data protection, aiming to protect both the computed function value and the underlying data. The work introduces a general framework for this joint protection requirement and derives corresponding redundancy bounds, adding another variant to the study of function-correcting codes.
@@ I-B Contributions and organization of the paper
  • We introduce $(\mathcal{P},t)$-encodings, called function-correcting partition codes, defined directly on a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{k}$. We show that any $(f,t)$-FCC is exactly an FCPC with respect to the domain partition induced by $f$, which makes FCPCs a natural generalization of FCCs.
  • We use the join of domain partitions to construct a single code that protects multiple functions simultaneously, and we specialize this to the linear functions via coset partition of the intersection of their kernels.
  • We define the notion of partition redundancy gain and partition rate gain to measure the bandwidth saved by using a single FCPC for multiple functions instead of constructing separate FCCs for each function. We derive general lower and upper bounds on the redundancy of such FCPCs and illustrate the achievable gains through examples.
  • We associate a partition graph to any given partition of $\mathbb{F}_{q}^{k}$, and show that the existence of an appropriately chosen clique in this graph allows the problem of determining the optimal redundancy to be reduced to analyzing only the vectors corresponding to the vertices of that clique. Then we show the existence of such cliques for three classes of partitions: coset partitions, Hamming-weight partitions, and support partitions.
@@
  • We introduce the notion of a block-preserving contraction for a partition, which helps reduce the problem of finding optimal redundancy for an FCPC. We further show that such a contraction exists for all weight-based partitions. We also give a condition for cosets under which the partition graph of the corresponding coset partition admits a block-preserving contraction.
@@
The rest of the paper is organized as follows. Section II recalls the basic notions on partitions, graphs, and function-correcting codes. In Section III, we introduce FCPCs, study their relation to FCCs, consider their use for protecting multiple functions, and define the partition gains together with general redundancy bounds, including the case of linear functions. Section IV introduces the partition graph and then focuses on coset partitions, weight-based partitions and support partition. Section V presents the notion of block-preserving contraction of partitions and establishes its connection to the optimal redundancy of the corresponding FCPC. Section VI discusses the notion of structural privacy inherent to FCPCs and illustrates the resulting ambiguity among functions inducing the same domain partition. Finally, Section VII concludes the paper.
@@ II Preliminaries
This section contains some basic definitions related to partitions [15], graphs [16] and function-correcting codes [1].
@@
Consider a function $f:\mathbb{F}_{q}^{k}\rightarrow S$. A systematic encoding $\mathcal{C}:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{k+r}$ is defined as an $(f,t)$-function correcting code (FCC) if, for any $u_{1},u_{2}\in\mathbb{F}_{2}^{k}$ such that $f(u_{1})\neq f(u_{2})$, the following condition holds:
@@
Let $\mathcal{D}\in\mathbb{N}^{M\times M}$. Then $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$ where $p_{i}\in\mathbb{F}_{q^{r}},~~i=1,2,\cdot,M$ is said to be an irregular-distance code or $\mathcal{D}$-code if there is an ordering of $P$ such that $d(p_{i},p_{j})\geq[\mathcal{D}]_{i,j}$ for all $i,j\in\{1,2,\ldots,M\}$. Further, $N(\mathcal{D})$ is defined as the smallest integer $r$ such that there exists a $\mathcal{D}$-code of length $r$. If $\mathcal{D}]_{i,j}=D$ for all $i,j\in\{1,2,\ldots,M\},i\neq j$, then $N(\mathcal{D})$ is denoted as $N(M,D)$.
@@ III Function-correcting partition codes
The optimal redundancy $r_{\mathcal{P}}(k,t)$ is defined as the minimum of $r$ for which there exists an $(\mathcal{P},t)$-encoding $\mathcal{C}:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$.
@@
Let $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{E}\}$ be a partition of $\mathbb{F}_{q}^{k}$ and let $u_{1},u_{2},\ldots,u_{M}\in\mathbb{F}_{q}^{k}$. The partition distance requirement matrix $\mathcal{D}_{\mathcal{P}}(t,u_{1},\ldots,u_{M})$ for an $(\mathcal{P},t)$-encoding is a $M\times M$ matrix with entries
@@
For $\mathcal{D}=\mathcal{D}_{\mathcal{P}}(t,u_{1},u_{2},\ldots,u_{q^{k}})$, if we have a $\mathcal{D}$-code $Z=\{z_{1},z_{2},\ldots,z_{q^{k}}\}$, then it can be used to construct an $(\mathcal{P},t)$-encoding $\mathcal{C}(u_{i})=(u_{i},z_{i})$ for all $i\in\{1,2,\ldots,q^{k}\}$.
@@
Let $\mathcal{P}$ and $\mathcal{Q}$ be two partitions of $\mathbb{F}_{q}^{k}$. If $\mathcal{Q}$ is a refinement of $\mathcal{P}$, then any $(\mathcal{Q},t)$-encoding is also an $(\mathcal{P},t)$-encoding, and
@@
Therefore, $\mathcal{C}$ is also an $(\mathcal{P},t)$-encoding, and the optimal redundancy of it is bounded as $r_{\mathcal{P}}(k,t)\leq r_{\mathcal{Q}}(k,t).$
An $(n,q^{k},2t+1)$ error-correcting code is an $(\mathcal{Q},t)$-encoding for the finest partition $\mathcal{Q}=\big\{\{x\}:x\in\mathbb{F}_{q}^{k}\big\}$. Since this finest partition is a refinement of any partition of $\mathbb{F}_{q}^{k}$, such a code is an $(\mathcal{P},t)$-encoding for every partition $\mathcal{P}$.
@@
A $(\mathcal{P},t)$-encoding defined above for a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{k}$, is an $(f,t)$-function correcting code for any function $f:\mathbb{F}_{q}^{k}\rightarrow S$ with domain partition $\mathcal{P}_{f}=\mathcal{P}$.
@@
Let $h:\mathbb{F}_{q}^{k}\to S$ be a function, and let $g:\mathrm{Im}(h)\to S^{\prime}$ be a function defined on the image of $h$. Define $f:\mathbb{F}_{q}^{k}\to S^{\prime}$ by $f(x)=g(h(x))$ for all $x\in\mathbb{F}_{q}^{k}.$ Then, for every integer $t\geq 0$, we have
@@
Let $\mathcal{P}_{h}$ and $\mathcal{Q}_{f}$ denote the domain partitions of $h$ and $f$, respectively. We have
$\mathcal{P}_{h}=\{P_{s}:s\in\mathrm{Im}(h)\},\quad\text{where}\ P_{s}=\{x\in\mathbb{F}_{q}^{k}:h(x)=s\},$
$\mathcal{Q}_{f}=\{Q_{s^{\prime}}:s^{\prime}\in\mathrm{Im}(f)\},\quad\text{where}\ Q_{s^{\prime}}=\{x\in\mathbb{F}_{q}^{k}:f(x)=s^{\prime}\}.$
By the definition of $f$, for any $x\in\mathbb{F}_{q}^{k}$, we have $f(x)=g(h(x))$. For a fixed $s^{\prime}\in\mathrm{Im}(f)$,
$Q_{s^{\prime}}=\{x\in\mathbb{F}_{q}^{k}:g(h(x))=s^{\prime}\}=\bigcup_{s\in g^{-1}(s^{\prime})}\{x\in\mathbb{F}_{q}^{k}:h(x)=s\}=\bigcup_{s\in g^{-1}(s^{\prime})}P_{s}.$
Thus each block $Q_{s^{\prime}}$ of $\mathcal{Q}_{f}$ is a union of blocks of $\mathcal{P}_{h}$. In other words, $\mathcal{P}_{h}$ is a refinement of $\mathcal{Q}_{f}$. Now using Lemma 2 we get
$r_{\mathcal{Q}_{f}}(k,t)\leq r_{\mathcal{P}_{h}}(k,t),$
and equivalently, $r_{f}(k,t)\leq r_{h}(k,t)$ as claimed.
Further, if $g$ is a bijection on $\mathrm{Im}(h)$, then for all $s^{\prime}\in\mathrm{Im}(f)$, we have $|g^{-1}(s^{\prime})|=1$ and
$Q_{s^{\prime}}=\bigcup_{s\in g^{-1}(s^{\prime})}P_{s}=P_{s}\quad\text{for some}\ s\in\mathrm{Im}(h).$
Also $|\mathrm{Im}(h)|=|\mathrm{Im}(f)|$, which means $\mathcal{P}_{h}=\mathcal{Q}_{f}.$ Therefore $r_{f}(k,t)=r_{\mathcal{Q}_{f}}(k,t)=r_{\mathcal{P}_{h}}(k,t)=r_{h}(k,t).$

### v2

# arXiv:2601.06450v2[cs.IT] 02 Mar 2026

@@ Abstract
We introduce function-correcting partition codes (FCPCs), which are a natural generalization of function-correcting codes (FCCs). An FCPC is defined directly on a partition of the message space, rather than on a specific target function. We show that any $t$-error correcting code for a function $f$ is exactly an FCPC with respect to the domain partition induced by $f$, which makes these codes a natural generalization of FCCs. We use the join of domain partitions to construct a single code that protects multiple functions simultaneously, with application for scenarios where a receiver needs to protect more than one function value or in a broadcast scenario where the different receivers need protection for different function values. We define the notions of partition gains to measure the bandwidth saved by using a single FCPC for multiple functions instead of constructing separate FCCs for each function. We derive general lower and upper bounds on the redundancy of such FCPCs and illustrate the achievable gains through examples. We specialize this concept of using single code for protecting multiple functions to linear functions via coset partition of the intersection of their kernels. We also present explicit FCPC constructions for locally bounded partitions and grouped weight partitions. Then, we associate a partition graph with any given partition of $\mathbb{F}_{q}^{k}$, and show that the existence of a suitable clique in this graph yields a set of representative information vectors that achieves the optimal redundancy. Using the existence of a full-size clique in the weight partition and support partition, we obtain lower and upper bounds on the optimal redundancy of FCPCs for these partitions. We introduce the notion of a block-preserving contraction for a partition, which helps reduce the problem size of finding optimal redundancy for an FCPC. We further show that such a contraction exists for all weight-based partitions. Finally, we observe that FCPCs naturally provide a form of partial privacy in the sense that only the domain partition of the function needs to be revealed to the transmitter.
@@ I Introduction
The theory of function-correcting codes (FCCs) has so far been developed primarily for the standard one-to-one communication setting, where a single receiver is interested in obtaining the value of a particular function of the transmitted message vector. In this work, we move beyond that setting and consider a more general and practically relevant scenario. A transmitter broadcasts a message vector over a shared and error-prone link to multiple receivers, and each receiver is interested in a different function or attribute of the same message vector. Since these receivers may require protection against errors for their respective functions, the natural question is whether one can design a single code that simultaneously protects all the desired functions from up to $t$ errors.
A straightforward approach is to encode the entire message vector using a standard error-correcting code (ECC), thereby enabling complete recovery at all receivers. This trivially protects all functions but may introduce significantly more redundancy than necessary. Given that FCCs are typically studied in systematic form, another equally straightforward approach is to take the redundancies required for the different FCCs corresponding to the individual functions, append all of them to the message vector, and transmit the resulting longer vector. In this way, each receiver can simply use the redundancy meant for its own function to decode its desired value. However, this solution also incurs redundancy equal to the sum of all the individual FCC redundancies, which may be unnecessarily large. If, instead, one can find a single unified FCC that protects all the functions with strictly smaller redundancy, it leads to substantial bandwidth savings compared to both of the above approaches: (1) using an ECC that protects the entire message vector, and (2) using the sum of the redundancies of the separate FCCs designed for each function individually.
The aim is to construct codes that perform this task as required. Instead of working with the function values directly, we define codes based on partitions of the domain and refer to them as function-correcting partition codes (FCPCs). The setup is illustrated in Fig. 1. Since every function naturally partitions its domain into preimage sets, these FCPCs can be used as FCCs. In fact, each function induces a natural partition of the domain into its preimage sets, and an FCC is exactly the FCPC defined with respect to this induced partition. This makes FCPCs a direct generalization of FCCs.
@@
Further, we refer to the savings, arising from using a single code for multiple functions instead of using separate FCCs for each function, as the partition redundancy gain and relative rate increment.
There is another benefit of using FCPCs in a network setting, as they provide a degree of partial privacy regarding the function being computed, which we refer to as function class privacy. Here, a function class consists of all functions on the domain that induce the same domain partition. In FCCs, both the receiver and the transmitter must know the exact function. In contrast, an FCPC requires the transmitter to know only the domain partition induced by the function that the receiver wishes to compute. Since many different functions can belong to the same function class and induce the same domain partition, the transmitter cannot infer the exact function from the partition alone. Further discussion of this notion is given in Section VI.
@@ I-A Related works
The notion of FCCs was formally introduced in [1], which laid the foundational framework for studying error correction when the objective is to reliably recover a function of the transmitted message rather than the message itself. In this work, the authors defined the key parameters of FCCs, established fundamental redundancy bounds, and investigated structural properties of such codes. They also provided explicit constructions for several important classes of functions, including locally binary functions, the Hamming weight function, and the Hamming weight distribution function. For some of these cases, the proposed constructions were shown to be optimal by achieving the corresponding lower bounds on redundancy. The initial development focused on communication over binary symmetric channels.
Subsequent research has extended this framework to more general channel models. In [2], FCCs were studied in the context of symbol-pair read channels over binary fields, and this line of work was further generalized to $b$-symbol read channels in [3]. In [4], the authors leveraged the fact that the kernel of a linear function forms a subspace of $\mathbb{F}_{q}^{k}$, and used the associated coset partition to construct FCCs for linear functions. They also derived a general lower bound on the redundancy of FCCs, which specializes to a classical lower bound for ECCs in the case of bijective functions, since FCCs reduce to ECCs in this setting. The tightness of this bound was established for certain parameters. The work [5] further refined the study of FCCs for the Hamming weight and Hamming weight distribution functions, providing improved bounds and constructions tailored to these specific functions.
@@
Another line of work studies FCCs under alternative distance measures and channel models. In [9, 10], FCCs are investigated in the Lee metric, where Plotkin-type bounds are derived for general functions and explicit constructions are provided for several function classes, with optimality established in certain cases. The work [11] considers FCCs with homogeneous distance, developing redundancy bounds and constructions adapted to this setting. In addition, Plotkin-type bounds for FCCs over $b$-symbol read channels are studied in [12].
More recently, a related direction was explored in [13], where the authors study FCCs with data protection, aiming to protect both the computed function value and the underlying data. The work introduces a general framework for this joint protection requirement and derives corresponding redundancy bounds, adding another variant to the study of FCCs.
@@ I-B Contributions and organization of the paper
  • We introduce $(\mathcal{P},t)$-encodings, called function-correcting partition codes, defined directly on a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{k}$. We show that any FCC is exactly an FCPC with respect to the domain partition induced by $f$, which makes FCPCs a natural generalization of FCCs.
  • We use the join of domain partitions to construct a single code that protects multiple functions simultaneously, and we specialize this to linear functions via coset partition of the intersection of their kernels.
  • We define the notions of partition redundancy gain and relative rate increment to measure the bandwidth saved by using a single FCPC for multiple functions instead of constructing separate FCCs for each function. We derive general lower and upper bounds on the redundancy of such FCPCs and illustrate the achievable gains through examples.
  • We present an explicit optimal FCPC construction for locally bounded partitions, obtained as a reformulation of the FCC construction for locally bounded functions introduced in [1]. We also propose FCPC constructions for grouped weight partitions by employing Gray codes and extending the approach of [5]. Furthermore, for the subclass of consecutive grouped weight partitions, we identify conditions under which the proposed construction achieves optimal redundancy.
  • We associate a partition graph with any given partition of $\mathbb{F}_{q}^{k}$, and show that the existence of an appropriately chosen clique in this graph allows the problem of determining the optimal redundancy to be reduced to analyzing only the vectors corresponding to the vertices of that clique. Then we show the existence of such cliques for three classes of partitions: coset partitions, Hamming-weight partitions, and support partitions.
@@
  • We introduce the notion of a block-preserving contraction for a partition, which helps reduce the problem size of finding optimal redundancy for an FCPC. We further show that such a contraction exists for all weight-based partitions. We also give a condition for cosets under which the partition graph of the corresponding coset partition admits a block-preserving contraction.
@@
The rest of the paper is organized as follows. Section II recalls the basic notions on partitions, graphs, and FCCs. In Section III, we introduce FCPCs, study their relation to FCCs, consider their use for protecting multiple functions, and define the partition gains together with general redundancy bounds, including the case of linear functions. Section IV introduces the partition graph and then focuses on coset partitions, weight-based partitions and support partition. Section V presents the notion of block-preserving contraction of partitions and establishes its connection to the optimal redundancy of the corresponding FCPC. Section VI discusses the notion of structural privacy inherent to FCPCs and illustrates the resulting ambiguity among functions inducing the same domain partition. Finally, Section VII concludes the paper.
@@ II Preliminaries
This section contains some basic definitions related to partitions [15, 16], graphs [17] and FCCs [1].
@@
Consider a function $f:\mathbb{F}_{q}^{k}\rightarrow S$. A systematic encoding $\mathcal{C}:\mathbb{F}_{q}^{k}\rightarrow\mathbb{F}_{q}^{k+r}$ is defined as an $(f,t)$-function correcting code (FCC) if, for any $u_{1},u_{2}\in\mathbb{F}_{q}^{k}$ such that $f(u_{1})\neq f(u_{2})$, the following condition holds:
@@
Let $\mathcal{D}\in\mathbb{N}^{M\times M}$. Then $\mathcal{P}=\{p_{1},p_{2},\ldots,p_{M}\}$, where $p_{i}\in\mathbb{F}_{q}^{r},~i=1,2,\ldots,M$, is said to be an irregular-distance code or $\mathcal{D}$-code if there is an ordering of $P$ such that $d(p_{i},p_{j})\geq[\mathcal{D}]_{i,j}$ for all $i,j\in\{1,2,\ldots,M\}$. Further, $N_{q}(\mathcal{D})$ is defined as the smallest integer $r$ such that there exists a $\mathcal{D}$-code of length $r$. If $\mathcal{D}]_{i,j}=D$ for all $i,j\in\{1,2,\ldots,M\},i\neq j$, then $N_{q}(\mathcal{D})$ is denoted as $N_{q}(M,D)$.
@@ III Function-correcting partition codes
The optimal redundancy $r_{\mathcal{P}}(k,t)$ is defined as the minimum of $r$ for which there exists a $(\mathcal{P},t)$-encoding $\mathcal{C}:\mathbb{F}_{2}^{k}\rightarrow\mathbb{F}_{2}^{k+r}$.
@@
Let $\mathcal{P}=\{P_{1},P_{2},\ldots,P_{E}\}$ be a partition of $\mathbb{F}_{q}^{k}$ and let $u_{1},u_{2},\ldots,u_{M}\in\mathbb{F}_{q}^{k}$. The partition distance requirement matrix $\mathcal{D}_{\mathcal{P}}(t,u_{1},\ldots,u_{M})$ for a $(\mathcal{P},t)$-encoding is a $M\times M$ matrix with entries
@@
For $\mathcal{D}=\mathcal{D}_{\mathcal{P}}(t,u_{1},u_{2},\ldots,u_{q^{k}})$, if we have a $\mathcal{D}$-code $Z=\{z_{1},z_{2},\ldots,z_{q^{k}}\}$, then it can be used to construct a $(\mathcal{P},t)$-encoding $\mathcal{C}(u_{i})=(u_{i},z_{i})$ for all $i\in\{1,2,\ldots,q^{k}\}$.
@@
Let $\mathcal{P}$ and $\mathcal{Q}$ be two partitions of $\mathbb{F}_{q}^{k}$. If $\mathcal{Q}$ is a refinement of $\mathcal{P}$, then any $(\mathcal{Q},t)$-encoding is also a $(\mathcal{P},t)$-encoding, and
@@
Therefore, $\mathcal{C}$ is also a $(\mathcal{P},t)$-encoding, and the optimal redundancy of it is bounded as $r_{\mathcal{P}}(k,t)\leq r_{\mathcal{Q}}(k,t).$
An $(n,q^{k},2t+1)$ ECC is a $(\mathcal{Q},t)$-encoding for the finest partition $\mathcal{Q}=\big\{\{x\}:x\in\mathbb{F}_{q}^{k}\big\}$. Since this finest partition is a refinement of any partition of $\mathbb{F}_{q}^{k}$, such a code is a $(\mathcal{P},t)$-encoding for every partition $\mathcal{P}$.
@@
A $(\mathcal{P},t)$-encoding defined above for a partition $\mathcal{P}$ of $\mathbb{F}_{q}^{k}$, is an $(f,t)$-FCC for any function $f:\mathbb{F}_{q}^{k}\rightarrow S$ with domain partition $\mathcal{P}_{f}=\mathcal{P}$.
@@
Let $h:\mathbb{F}_{q}^{k}\to S$ be a function, and let $g:\mathrm{Im}(h)\to S^{\prime}$ be a function defined on the image of $h$. Consider $f:\mathbb{F}_{q}^{k}\to S^{\prime}$ defined by $f(x)=g(h(x))$ for all $x\in\mathbb{F}_{q}^{k}.$ Then, for every integer $t\geq 0$, we have
@@
Let $\mathcal{P}_{h}$ and $\mathcal{P}_{f}$ denote the domain partitions of $h$ and $f$, respectively. If $x,y\in\mathbb{F}_{q}^{k}$ are in the same block of $\mathcal{P}_{h}$, then $h(x)=h(y)$, so $f(x)=g(h(x))=g(h(y))=f(y)$, which means $x$ and $y$ are also in the same block of $\mathcal{P}_{f}$. Therefore, $\mathcal{P}_{h}$ is a refinement of $\mathcal{P}_{f}$, so by Lemma 2 we get $r_{\mathcal{P}_{f}}(k,t)\leq r_{\mathcal{P}_{h}}(k,t),$ and equivalently, $r_{f}(k,t)\leq r_{h}(k,t)$ as claimed. Furthermore, if $g$ is bijective on $\mathrm{Im}(h)$ then we can also write $h=g^{-1}\circ f$, so the partitions $\mathcal{P}_{h}$ and $\mathcal{P}_{f}$ are refinements of each other, and therefore equal as partitions. It follows that $r_{f}(k,t)=r_{\mathcal{P}_{f}}(k,t)=r_{\mathcal{P}_{h}}(k,t)=r_{h}(k,t).$